### v1

# Uncertainty-Aware 4D Gaussian Splatting for Monocular Occluded Human Rendering

Abstract
High-fidelity rendering of dynamic humans from monocular videos typically degrades catastrophically under occlusions. Existing solutions incorporate external priors—either hallucinating missing content via generative models, which induces severe temporal flickering, or imposing rigid geometric heuristics that fail to capture diverse appearances. To this end, we reformulate the task as a Maximum A Posteriori estimation problem under heteroscedastic observation noise. In this paper, we propose U-4DGS, a framework integrating a Probabilistic Deformation Network and a Double Rasterization pipeline. This architecture renders pixel-aligned uncertainty maps that act as an adaptive gradient modulator, automatically attenuating artifacts from unreliable observations. Furthermore, to prevent geometric drift in regions lacking reliable visual cues, we enforce Confidence-Aware Regularizations, which leverage the learned uncertainty to selectively propagate spatial-temporal validity. Extensive experiments on ZJU-MoCap and OcMotion demonstrate that U-4DGS achieves SOTA rendering fidelity and robustness.
1 Introduction
High-fidelity reconstruction of dynamic humans from monocular videos stands as a foundational technology for immersive applications, spanning telepresence, metaverse interaction, sports analysis, and virtual reality [tian2023recovering, yang2025innovative]. Recent years have witnessed remarkable progress in this domain, enabling the synthesis of photorealistic digital avatars from monocular inputs [weng2022humannerf, wen2024gomavatar]. However, these successes hinge on the idealized assumption of full visibility and isolation from the background. This assumption rarely holds in unconstrained in-the-wild scenarios, where subjects inevitably interact with their surroundings, such as sitting on chairs, navigating behind obstacles, or being partially obscured by environmental objects. When standard pipelines are applied to occluded scenarios, their performance degrades catastrophically [xiang2023rendering, sun2024occfusion]. As illustrated in Fig. 1(e), the model erroneously interprets the occluder as part of the human geometry, manifesting as incomplete geometry or severe texture artifacts. Consequently, enabling robust human rendering from occluded observations is imperative to bridge the gap between laboratory prototypes and real-world applications.
To address these specific challenges, pioneering NeRF-based approaches, such as OccNeRF [xiang2023rendering] and Wild2Avatar [xiang2025rendering], explore the task of rendering humans from object-occluded monocular videos. By incorporating geometry priors or decoupling scene components, these methods demonstrate the feasibility of recovering occluded surfaces. However, despite their visual fidelity, they inherit the intrinsic computational bottlenecks of implicit neural representations. Specifically, the reliance on expensive volumetric rendering incurs prohibitive training and inference costs, effectively precluding their deployment in real-time applications.
Recently, the advent of 3D Gaussian Splatting (3DGS) [kerbl20233d] shifted the paradigm, offering real-time rendering capabilities [fei20243d, feng2025flashgs]. To adapt 3DGS for occluded human rendering, concurrent works incorporate external priors to compensate for missing information. Methods relying on geometric heuristics enforce rigid rules, such as left-right symmetry [jiang2025symgaussian] or feature aggregation [ye2025occgaussian]. However, these hand-crafted spatial priors typically lack flexibility in modeling diverse human appearances. As shown in Fig. 1(f), the symmetry prior fails to recover the chest logo, erroneously copying unadorned features from the opposite side. Alternatively, methods leveraging generative priors utilize 2D diffusion models to hallucinate invisible regions [sun2024occfusion, lee2024guess]. Yet, the inherent stochasticity of frame-by-frame generation often induces severe temporal flickering. This is evident in Fig. 1(g), where the shirt color drifts inconsistently across frames, violating physical consistency. Crucially, these methods rely on imperfect external priors to infer missing content, neglecting the intrinsic reliability of the input data, where valid cues are embedded across the temporal sequence.
To bridge this gap, our insight stems from the video’s intrinsic temporal redundancy: since occlusions are transient, parts obscured in one frame are likely visible in others. This motivates adopting a 4D representation to consolidate fragmented observations into a unified canonical space. However, naive 4D Gaussian Splatting faces a supervision dilemma [wu20244d, li2024st]. Indiscriminately treating all pixels as valid constraints forces the deformation field to establish erroneous correspondences, compelling the human geometry to adhere to the occluder’s surface texture. Conversely, strictly limiting supervision to visible regions leaves the occluded parts mathematically unconstrained; without gradient feedback, the geometry in these blind spots fails to densify or maintain coherence, resulting in incomplete structures. Thus, the reconstruction process demands a dynamic discrimination mechanism to assess the validity of the observation, differentiating between visible human details and misleading observational noise (i.e., occluded regions).
To this end, we reformulate the task as a Maximum A Posteriori (MAP) estimation problem under heteroscedastic observation noise. Specifically, instead of assuming constant variance across all observations, we model the likelihood of each pixel as a Laplacian distribution with a learnable scale. This formulation naturally derives a probabilistic objective where the predicted scale—representing the aleatoric uncertainty—emerges as an inverse weighting term in the loss function. Consequently, this uncertainty functions as an adaptive gradient modulator: low uncertainty permits RGB supervision to dominate the optimization for fine-grained refinement, while high uncertainty automatically attenuates the gradients from unreliable observations. As visualized in Fig. 1(b), the network explicitly captures this reliability distribution, autonomously assigning high uncertainty (bright regions) to occluded areas.
Building upon this formulation, we present U-4DGS, a robust uncertainty-aware framework for monocular occluded human rendering. To parameterize the heteroscedastic noise model, we propose a Probabilistic Deformation Network. Beyond standard geometric warping, this module explicitly predicts the per-primitive aleatoric uncertainty, serving as the confidence metric for our MAP objective. To enable the adaptive gradient modulation, we introduce a Double Rasterization pipeline. By rendering pixel-aligned uncertainty maps alongside color, this mechanism enforces the inverse weighting scheme, ensuring that gradients from occluded regions are effectively attenuated. Finally, to prevent geometric drift in areas lacking visual cues (i.e., occluded regions), we design Confidence-Aware Regularizations. These constraints leverage the learned uncertainty to selectively propagate validity from confident regions to unreliable areas, ensuring coherent completion. Consequently, as illustrated in Fig. 1(a), U-4DGS achieves high-fidelity rendering even under severe occlusion while maintaining training efficiency.
In summary, our main contributions are as follows:
We reformulate occluded human rendering as a MAP estimation problem. By modeling heteroscedastic observation noise, we equip the optimization with a mechanism to discriminate between valid supervision and occlusion artifacts.
We propose U-4DGS, a framework integrating a Probabilistic Deformation Network with a Double Rasterization pipeline. This design renders pixel-aligned uncertainty maps that act as adaptive gradient modulators, effectively shielding the canonical geometry from corruption.
We introduce Confidence-Aware Regularizations to prevent geometric drift in regions lacking reliable visual cues. These constraints leverage learned uncertainty to enforce spatial-temporal consistency, ensuring coherent completion.
We demonstrate that U-4DGS achieves state-of-the-art rendering fidelity on ZJU-Mocap and OcMotion datasets, successfully recovering clean avatars from severe occlusions.
2 Related Works
3D Human Avatar Reconstruction. Traditional human reconstruction methods typically rely on dense camera arrays or depth sensors, limiting their application in in-the-wild scenarios [collet2015high, su2020robustfusion, huang2020arch, yu2021function4d]. To address this, NeRF-based methods have enabled high-fidelity reconstruction from monocular inputs by conditioning radiance fields on SMPL priors [yu2023monohuman, te2022neural, li2024ghunerf]. However, the reliance on volumetric ray-marching incurs prohibitive computational costs, precluding real-time applicability. Recently, 3D Gaussian Splatting (3DGS) [kerbl20233d] has introduced a paradigm shift with its real-time rasterization capabilities. State-of-the-art methods [qian20243dgs, kocabas2024hugs, hu2024gauhuman, lei2024gart, shao2024splattingavatar, zhao2025surfel] adapt this representation to dynamic humans by representing the body in a canonical space and transforming it via Linear Blend Skinning. Critically, while performant under full visibility, these standard pipelines implicitly assume valid supervision for all pixels. Thus, they lack mechanisms to distinguish the human subject from occlusions, leading to catastrophic degradation where occluder artifacts are erroneously fused into the avatar geometry.
Occlusion-Aware Human Rendering. Existing methodologies generally fall into three paradigms. Early strategies focused on scene decoupling, employing visibility priors to isolate the subject from obstacles [xiang2025rendering, xiang2023rendering]; however, these implicit frameworks often suffer from high computational overhead. To synthesize unobserved content, recent approaches leverage generative priors via pre-trained 2D diffusion models [sun2024occfusion, lee2024guess, fan2026inpainthuman]. Yet, the inherent stochasticity of such processes frequently induces identity drift and temporal flickering. Alternatively, other 3DGS-based methods employ geometric heuristics such as symmetry constraints [ye2025occgaussian, jiang2025symgaussian], which often falter under asymmetric clothing or motions. Distinctively, U-4DGS eschews unstable hallucinations and rigid heuristics. Instead, we formulate the reconstruction as an MAP estimation problem. By strictly aggregating valid temporal information via an adaptive gradient modulation mechanism, we ensure reconstruction that is both physically consistent and faithful to the subject.
Dynamic Reconstruction and Uncertainty Modeling. Recent strides in dynamic Gaussian Splatting have achieved high-fidelity 4D reconstruction via deformation fields [yang2024deformable, wu20244d, luiten2024dynamic, bae2024per, zhang2025mega]. However, these methods predicate their success on a deterministic correspondence between canonical and observation spaces. While valid for clean sequences, this premise disintegrates under occlusion, where the absence of valid motion cues causes the solver to overfit to occluders [guo2025uncertainty]. To address this, uncertainty modeling has been explored to enhance robustness against sparse views or transient noise [sunderhauf2023density, goli2024bayes, ren2024nerf]. Notably, concurrent work like USPLAT4D [guo2025uncertainty] employs an uncertainty-aware motion graph for general dynamic scenes. Yet, treating uncertainty solely as a smoothing weight or edge attribute proves insufficient for articulated human reconstruction, where structural integrity is paramount. In contrast, U-4DGS elevates uncertainty to a central arbitration role. Beyond simple loss attenuation, we leverage learned confidence to actively modulate optimization dynamics: shielding canonical geometry from misleading gradients while simultaneously triggering confidence-aware regularizations to enforce physical plausibility within regions lacking reliable visual cues.
3 Preliminaries
3.1 3D Gaussian Splatting
3D Gaussian Splatting (3DGS) [kerbl20233d] explicitly models the scene using a set of anisotropic 3D Gaussians. Each primitive is parameterized by a mean , covariance , opacity , and view-dependent color (via Spherical Harmonics). To enforce positive semi-definiteness during differentiable optimization, the covariance is decomposed into a rotation quaternion and a scaling vector :
where and denote the rotation and scaling matrices, respectively.
For rendering, Gaussians are projected onto the image plane to compute the pixel color via differentiable front-to-back -blending of sorted primitives:
where represents the probability density of the projected 2D Gaussian at pixel , and denotes the accumulated transmittance.
3.2 Articulated Human Gaussians
Following recent works [hu2024gauhuman, qian20243dgs], we leverage the SMPL model [loper2015smpl] to establish a canonical coordinate system. SMPL provides a skinned vertex mesh parametrized by pose and shape . To model dynamic humans, we instantiate 3D Gaussians within this canonical T-pose space. For a specific frame with pose , any canonical point is transformed to the observation space via Linear Blend Skinning:
where denotes the LBS function, represents the joint locations, and contains the blend weights. In U-4DGS, we bind Gaussian primitives to their nearest SMPL vertices. This strategy ensures that coarse body movements are explicitly driven by skeletal articulation, while a learned neural field subsequently models residual non-rigid deformations (e.g., clothing dynamics).
3.3 Problem Formulation: A Probabilistic Perspective
Given a monocular video sequence capturing a dynamic human, our objective is to reconstruct a high-fidelity human avatar. We assume access to camera parameters and coarse SMPL tracking. From a probabilistic standpoint, standard 3DGS optimization can be viewed as a Maximum Likelihood Estimation (MLE) problem. Minimizing the standard photometric loss (e.g., in Eq. 2) implicitly assumes that the observation noise across all pixels follows a fixed, homoscedastic distribution (i.e., constant variance). While effective for clean sequences, this assumption disintegrates under environmental occlusion. In such scenarios, the observation noise is inherently heteroscedastic: visible regions exhibit low noise, whereas occluded regions suffer from gross corruption. Applying a homoscedastic MLE objective to such data forces the model to overfit these outliers, compelling the human geometry to adhere to the occluder’s surface erroneously.
To resolve this, we move beyond deterministic fitting and reformulate the reconstruction as a Maximum A Posteriori (MAP) estimation problem. Our goal is to jointly estimate the canonical parameters and the per-pixel aleatoric uncertainty (representing the noise scale). This formulation allows us to dynamically down-weight unreliable observations while incorporating structural priors to constrain the solution in regions lacking reliable visual cues. We detail this uncertainty-aware framework in Sec. 4.
4 Methodology
4.1 Overview
Building upon the probabilistic formulation derived in Sec. 3, we propose U-4DGS, a framework designed to realize the MAP estimation for robust occluded human rendering. U-4DGS explicitly models the heteroscedastic nature of observation noise, enabling an autonomous arbitration mechanism between valid supervision and occlusion artifacts. As illustrated in Fig. 2, the pipeline consists of three integral components: (1) Probabilistic Deformation (Sec. 4.2): We introduce a network that serves a dual role: it models the temporal geometry evolution via deformation offsets and parameterizes the input-dependent aleatoric uncertainty for each Gaussian, explicitly quantifying the confidence of the temporal correspondence. (2) Double Rasterization (Sec. 4.3): To propagate this 3D confidence to the 2D observation space, we employ a dual-branch rasterizer. This module renders the deformed geometry into a photometric image while simultaneously accumulating per-primitive uncertainty into a pixel-aligned Uncertainty Map . (3) Optimization via MAP Estimation (Sec. 4.4): These outputs drive our MAP-based objective. Here, functions as an adaptive gradient modulator within the probabilistic photometric loss, attenuating gradients from unreliable observations. Simultaneously, to prevent geometric drift in regions lacking reliable visual cues, we enforce Confidence-Aware Regularizations that leverage the learned uncertainty to impose spatial-temporal priors.
4.2 Probabilistic Deformation Network
To instantiate the MAP framework, we require a mechanism that not only recovers time-varying geometry but also quantifies the reliability of each correspondence. Unlike standard deterministic fields [yang2024deformable, hu2024gauhuman] that indiscriminately fit all observations, we propose a Probabilistic Deformation Network . This module serves a dual purpose: (1) it models high-frequency non-rigid residuals beyond the SMPL prior; (2) it explicitly predicts the aleatoric uncertainty , providing the input-dependent noise scale required for our probabilistic objective.
We parameterize as a Multi-Layer Perceptron (MLP). To capture high-frequency details in both spatial and temporal domains, we apply sinusoidal positional encoding to both the input coordinates and time. The network conditions the canonical Gaussians on these embeddings and the 3D skeleton pose to predict geometry offsets and uncertainty:
where denotes the stop-gradient operator. The offsets explicitly model the time-varying geometry, enabling the fusion of features across frames. For the uncertainty , we employ a Softplus activation to ensure strictly positive values (). We then apply the predicted offsets to the canonical Gaussians to obtain the deformed state:
These deformed primitives are subsequently transformed to the observation space via LBS (Eq. 3).
Crucially, since is predicted in the canonical space, it is defined as an intrinsic property of the Gaussian primitive. It remains bound to the surface during the LBS transformation, invariant to the global body movement. This design allows the network to learn specific spatial-temporal failure modes, such as assigning consistently high uncertainty to body parts undergoing rapid deformation or frequent self-occlusion, regardless of their global position in the camera view.
4.3 Double Rasterization
To bridge the gap between the 3D aleatoric uncertainty and the 2D observation space required for our MAP objective, we propose a Double Rasterization scheme. Since occlusion is an inherently view-dependent phenomenon, the reliability of a pixel is determined by the accumulated uncertainty along its corresponding optical ray. Therefore, we perform two pixel-aligned rasterization passes in parallel, generating both the photometric appearance and the uncertainty map .
The first pass renders the predicted color image using the standard splatting formulation (Eq. 2). Crucially, this pass utilizes the deformed geometry derived from the Probabilistic Deformation Network, ensuring that the rendering faithfully reflects the current temporal dynamics.
The second pass renders the uncertainty map . To ensure strict geometric consistency between the rendered appearance and its associated confidence, we accumulate the uncertainty using the same -blending weights used for color:
where and are identical to those computed in the color pass. This formulation has a clear physical interpretation: represents the expected aleatoric uncertainty of the visible surface at pixel . By compositing with the opacity , the resulting map accurately reflects the confidence of the foremost visible surface, automatically disregarding occluded primitives hidden behind. Consequently, serves as the spatially varying noise scale, functioning as the adaptive gradient modulator in the subsequent optimization.
4.4 Optimization via MAP Estimation
To reconstruct high-fidelity humans from monocular videos, we jointly optimize the canonical Gaussian parameters and the deformation network weights . Following the MAP formulation introduced in Sec. 3, our objective function maximizes the posterior . This decomposes the optimization into two synergistic components: a Likelihood term that models data fidelity under heteroscedastic noise, and a Prior term that enforces physical constraints via Confidence-Aware Regularizations.
Uncertainty-Weighted Photometric Loss (). The primary challenge in occluded rendering is the mismatch between the reconstructed geometry and occluder-corrupted observations. Standard objectives (e.g., L1 loss) assume homoscedastic noise, treating all pixels equally and thus forcing the model to fit occlusion artifacts. To resolve this, we model the observation likelihood by assuming the pixel-wise residual follows a Laplacian distribution, where the location parameter is the predicted color and the scale parameter is the predicted uncertainty . The probability density function is given by . Minimizing the negative log-likelihood (NLL) of this distribution yields our uncertainty-weighted objective:
where ensures numerical stability. This objective functions as an adaptive gradient modulator: in visible regions, the solver minimizes the numerator ( error) and reduces ; in occluded regions, it autonomously increases to attenuate the gradient magnitude. This mechanism effectively shields the canonical geometry from misleading supervision.
Confidence-Aware Regularizations (). While the uncertainty-weighted likelihood effectively attenuates gradients from artifacts, it leaves regions lacking reliable visual cues mathematically unconstrained. Without valid supervision, the human Gaussians in these blind spots are prone to the ill-posed geometric drift mentioned in Sec. 1. To resolve this, we impose the prior via Confidence-Aware Regularizations. These constraints utilize the learned uncertainty to selectively enforce physical plausibility: applying strong regularization only where observations are unreliable, while allowing data-driven deformation in confident regions.
(1) Confidence-Guided Spatial Consistency: To maintain geometric integrity, we enforce local rigidity weighted by uncertainty. We construct a KNN graph in the canonical space. For each Gaussian and its neighbor , we constrain their deformation attributes to be consistent:
Here, weights neighbors by canonical distance. The term acts as an attention mechanism: When a Gaussian is occluded ( is high), the gradients vanish, and penalty dominates, forcing the Gaussian to synchronize with its neighbors. Crucially, since visible neighbors are firmly anchored by the photometric loss, this constraint effectively propagates geometric validity from the confident boundary inward to the occluded regions.
(2) Uncertainty-Weighted Temporal Inertia: To prevent jittering in the absence of visual cues, we enforce a smoothness prior on the trajectory of uncertain Gaussians. Considering that we train on monocular video sequences with a frame interval , we formulate this constraint as a second-order difference over the sampled timestamps:
where denotes the set of deformation parameters, and represents the set of valid frames with available temporal neighbors. This term acts as an inertial prior. For high-uncertainty Gaussians, it penalizes acceleration, effectively interpolating their dynamics based on the trajectory established by confident frames, ensuring smooth motion transition across occlusion gaps.
Note that in both and , we apply a stop-gradient operator to the uncertainty weights . This ensures that serves purely as an attention mechanism for the regularization strength, preventing the optimizer from trivially reducing to minimize the regularization penalty.
Total Loss. The final objective function consolidates the MAP formulation, combining the uncertainty-weighted likelihood (data fidelity), the confidence-aware priors (regularization), and auxiliary perceptual constraints:
where ensures perceptual quality using standard metrics [wang2004image, zhang2018unreasonable]. Additionally, imposes a global silhouette constraint, aligning the accumulated opacity with the projected SMPL mask to guide coarse geometry. The hyperparameters balance the contribution of each term.
[…]
| Standard Human Rendering | |||||||||||||
| HumanNeRF [weng2022humannerf] | - | 20.67 | 0.9509 | - | - | - | - | ||||||
| GaussianAvatar [hu2024gaussianavatar] | - | 18.01 | 0.9512 | 60.33 | - | - | - | ||||||
| GauHuman [hu2024gauhuman] | - | 21.55 | 0.9430 | 55.88 | 15.09 | 0.8525 | 107.1 | ||||||
| Occlusion-Aware Approaches | |||||||||||||
| OccNeRF [xiang2023rendering] | Scene Decoupling | 22.40 | 0.9562 | 43.01 | 15.71 | 0.8523 | 82.90 | ||||||
| Wild2Avatar [xiang2025rendering] | - | - | - | 14.09 | 0.8484 | 93.31 | |||||||
| OccGaussian [ye2025occgaussian] | Geometric Heuristics | 23.29 | 0.9482 | 41.93 | - | - | - | ||||||
| SymGaussian [jiang2025symgaussian] | 23.22 | 0.9535 | 39.02 | - | - | - | |||||||
| GTU [lee2024guess] | Generative Priors | 22.89 | 0.9503 | 40.78 | 15.83 | 0.8437 | 83.46 | ||||||
| OccFusion [sun2024occfusion] | 23.96 | 0.9548 | 32.34 | 18.28 | 0.8875 | 82.42 | |||||||
| U-4DGS (Ours) | 24.62 | 0.9606 | 31.72 | 20.11 | 0.9030 | 79.17 | |||||||
5 Experiments
5.1 Experimental Setup
Datasets. We evaluate U-4DGS on two standard benchmarks representing both synthetic and real-world occlusion scenarios. First, we utilize the ZJU-MoCap dataset [peng2021neural], which contains multi-view sequences of 6 subjects. Following the protocol of OccNeRF [xiang2023rendering], we simulate occlusions by masking the central 50% of the human region in the first 80% of the frames. We adopt a monocular setting for training, using only Camera 1 with 100 frames subsampled at an interval of 5, while the remaining 22 cameras are reserved for novel view evaluation. Second, to assess robustness in natural environments, we employ the OcMotion dataset [huang2022occluded], which features varying degrees of real-world human-object interactions. Consistent with prior works [xiang2025rendering, sun2024occfusion], we select 6 representative sequences and train on sparse sub-sequences of 50 frames to challenge the model’s capability under limited observations.
Baselines. We benchmark our method against comprehensive state-of-the-art approaches, categorized into two distinct groups: (1) Standard Human Rendering. To demonstrate the impact of occlusion on conventional pipelines, we evaluate HumanNeRF [weng2022humannerf], GaussianAvatar [hu2024gaussianavatar], and GauHuman [hu2024gauhuman]. These methods are trained directly on the occluded sequences without specific handling mechanisms. (2) Occlusion-Aware Approaches. We compare against representative methods covering three mainstream technical paradigms: a) Scene Decoupling: OccNeRF [xiang2023rendering] and Wild2Avatar [xiang2025rendering], which utilize implicit fields or layer-wise parameterization to separate the human subject from environmental obstacles. b) Geometric Heuristics: OccGaussian [ye2025occgaussian] and SymGaussian [jiang2025symgaussian], which rely on rigid priors such as feature aggregation or left-right symmetry to infer missing geometry. c) Generative Priors: OccFusion [sun2024occfusion] and GTU [lee2024guess], which leverage 2D diffusion models to hallucinate the unobserved regions. For a fair comparison, all methods are evaluated using identical segmentation masks and SMPL pose priors.
Implementation Details: Our framework is implemented in PyTorch [paszke2019pytorch] and trained on a single NVIDIA A100 PCIe GPU. The Probabilistic Deformation Network is instantiated as an 8-layer MLP (hidden dimension 256) with a skip connection at the 4-th layer. To capture high-frequency spatial-temporal dynamics, we apply sinusoidal positional encoding to input coordinates and time with frequency bands and , respectively.
We optimize the model for 10k iterations using the Adam optimizer [kingma2014adam]. Learning rates for canonical Gaussian attributes follow the schedule of GauHuman [hu2024gauhuman], while the deformation network utilizes a rate of . Crucially, to preclude trivial solutions (where the model simply predicts infinite uncertainty to minimize the NLL), we adopt a progressive training strategy. For the initial 2k iterations, we deactivate the uncertainty branch and supervise geometry solely via standard loss. Subsequently, we activate the uncertainty head and switch to the MAP objective (Eq. 7) to jointly optimize for reconstruction fidelity and reliability. The regularization weights are set as: , , and .
5.2 Comparisons with State-of-the-Arts
Quantitative Results. Tab. 4.4 reports the performance on ZJU-MoCap and OcMotion benchmarks. On the synthetic ZJU-MoCap dataset, U-4DGS sets a new SOTA, achieving a PSNR of 24.62 dB. Notably, it surpasses the leading generative baseline, OccFusion (23.96 dB), validating that our uncertainty-guided aggregation yields superior fidelity without relying on potentially unstable diffusion priors. Crucially, on the real-world OcMotion dataset, our framework demonstrates exceptional robustness, outperforming the runner-up by a substantial margin of 1.83 dB. Unlike standard rendering or scene decoupling methods that degrade under occlusions, U-4DGS effectively leverages Confidence-Aware Regularizations to shield the canonical geometry, proving that modeling heteroscedastic noise is decisive for handling unconstrained interactions.
Qualitative Results. Fig. 3 visualizes the rendering quality. Standard methods like GauHuman (GH) [hu2024gauhuman] lack occlusion awareness, erroneously fusing environmental obstacles (e.g., bars or boxes) directly onto the human body. While generative approaches like OccFusion (OF) [sun2024occfusion] can remove occluders, they suffer from inherent stochasticity, often yielding over-smoothed textures or hallucinations inconsistent with the subject’s identity. In contrast, U-4DGS achieves superior fidelity. By functioning as an adaptive gradient modulator, our uncertainty mechanism effectively attenuates gradients from artifacts. By strictly aggregating valid temporal information, our method faithfully restores clean, sharp, and physically consistent avatars, even under severe occlusions.
5.3 Ablation and Analysis
Ablation Studies. We validate the efficacy of each component by incrementally integrating it into the baseline to instantiate our complete MAP framework. Quantitative metrics on ZJU-MoCap are reported in Tab. 2, and visual comparisons are presented in Fig. 4. (1) Exp. A (Baseline): We establish a deterministic baseline OccGauHuman, structurally inheriting the pipeline from GauHuman [hu2024gauhuman]. It employs a standard deformation field optimized via loss, effectively assuming homoscedastic observation noise. As shown in Fig. 4 (Exp. A), limited by this rigid deterministic fitting, the model fails to distinguish between the human body and environmental obstacles. Consequently, it indiscriminately minimizes geometric error, leading to severe artifact adhesion at occlusion boundaries. (2) Exp. B (+ Uncertainty Modeling): Replacing the deterministic field with our Probabilistic Deformation Network and enabling Double Rasterization yields a substantial quality leap. This step introduces the Likelihood term of our MAP objective. By functioning as an adaptive gradient modulator, the learned uncertainty effectively attenuates gradients from unreliable observations, successfully removing occlusion artifacts. However, without regularization priors, the geometry in regions lacking reliable visual cues remains noisy and ill-defined (see Exp. B), as the probabilistic loss solely down-weights error without explicitly guiding completion. (3) Exp. C (+ ): The introduction of Confidence-Guided Spatial Consistency imposes the first Prior term. By leveraging uncertainty to selectively propagate spatial validity from confident neighbors, this constraint effectively regularizes the noisy geometry, restoring a coherent and complete body shape. (4) Exp. D (+ ): Finally, incorporating Uncertainty-Weighted Temporal Inertia completes the U-4DGS framework. This temporal prior ensures trajectory smoothness and further refines surface details, achieving the most faithful reconstruction with high physical plausibility.
[…]
| A | Baseline | 22.85 | 0.9490 | 45.12 |
| B | + Uncertainty Modeling | 24.15 | 0.9575 | 35.20 |
| C | + Spatial Constraint () | 24.48 | 0.9592 | 32.50 |
| D | + Temporal Constraint () | 24.62 | 0.9606 | 31.72 |
Visualization of Uncertainty Maps. To validate the efficacy of our heteroscedastic noise modeling, we visualize the rendered uncertainty map in Fig. 5. In our MAP framework, represents the estimated scale of observation noise, tasked with autonomously distinguishing between valid signals and occlusion artifacts. As evident in Fig. 5, the network assigns low uncertainty (depicted in dark purple) to visible body parts, allowing the photometric supervision to drive the optimization in these trusted regions. Crucially, for anatomical regions that are occluded in the current frame but belong to the canonical topology, the network autonomously predicts significantly high uncertainty (depicted in bright yellow). This visualization empirically confirms that functions as an adaptive gradient modulator. By automatically increasing the denominator in the NLL objective (Eq. 7), it effectively attenuates gradients in these regions lacking reliable visual cues, preventing the canonical geometry from adhering to environmental occlusions.
Temporal Stability Analysis. To validate the temporal coherence of our framework, we conduct a longitudinal comparison against the generative baseline GTU [lee2024guess] across a long sequence (Fig. 6). As observed in the top row, GTU suffers from severe identity drift due to the inherent stochasticity of its diffusion priors. Since the method hallucinates missing regions frame-by-frame without strict temporal coupling, the synthesized textures fluctuate drastically. This instability becomes particularly pronounced in later frames (e.g., to ), where the clothing consistency collapses, degenerating into severe dark artifacts and incoherent noise on the subject’s back. In contrast, U-4DGS achieves rigorous stability. By anchoring appearance in a unified canonical space and enforcing Uncertainty-Weighted Temporal Inertia, our method effectively constrains the optimization trajectory. This ensures that the reconstructed human maintains physical consistency and sharp details throughout the sequence, eliminating the texture flickering observed in generative approaches.
6 Conclusion
In this paper, we present U-4DGS, a robust framework that reformulates monocular human rendering as an MAP estimation problem to handle severe environmental occlusion. Departing from paradigms that rely on stochastic hallucinations or rigid priors, we explicitly model the heteroscedastic nature of observation noise. By integrating a Probabilistic Deformation Network with a Double Rasterization pipeline, our method establishes an adaptive gradient modulator, enabling the optimization to autonomously distinguish between valid geometric cues and occlusion artifacts. Furthermore, we leverage the learned uncertainty to drive Confidence-Aware Regularizations, preventing geometric drift and ensuring structural integrity in regions lacking reliable visual cues. Extensive evaluations on ZJU-MoCap and OcMotion demonstrate that U-4DGS significantly outperforms state-of-the-art approaches, achieving superior robustness and photorealism.

### v2

@@
 Uncertainty-Aware 4D Gaussian Splatting for Monocular Occluded Human Rendering
-Abstract
+Abstract.
 High-fidelity rendering of dynamic humans from monocular videos typically degrades catastrophically under occlusions.
@@
 To this end, we reformulate the task as a Maximum A Posteriori estimation problem under heteroscedastic observation noise.
-In this paper, we propose U-4DGS, a framework integrating a Probabilistic Deformation Network and a Double Rasterization pipeline.
+In this paper, we propose U-4DGS, a framework integrating a Probabilistic Deformation Network and a Joint Rasterization pipeline.
 This architecture renders pixel-aligned uncertainty maps that act as an adaptive gradient modulator, automatically attenuating artifacts from unreliable observations.
 Furthermore, to prevent geometric drift in regions lacking reliable visual cues, we enforce Confidence-Aware Regularizations, which leverage the learned uncertainty to selectively propagate spatial-temporal validity.
-Extensive experiments on ZJU-MoCap and OcMotion demonstrate that U-4DGS achieves SOTA rendering fidelity and robustness.
-1 Introduction
-High-fidelity reconstruction of dynamic humans from monocular videos stands as a foundational technology for immersive applications, spanning telepresence, metaverse interaction, sports analysis, and virtual reality [tian2023recovering, yang2025innovative].
-Recent years have witnessed remarkable progress in this domain, enabling the synthesis of photorealistic digital avatars from monocular inputs [weng2022humannerf, wen2024gomavatar].
+Extensive experiments on the ZJU-MoCap and OcMotion datasets demonstrate that U-4DGS achieves state-of-the-art rendering fidelity and robustness.
+1. Introduction
+High-fidelity reconstruction of dynamic humans from monocular videos stands as a foundational technology for immersive applications, spanning telepresence, metaverse interaction, sports analysis, and virtual reality (Tian et al., 2023; Yang et al., 2025).
+Recent years have witnessed remarkable progress in this domain, enabling the synthesis of photorealistic digital avatars from monocular inputs (Weng et al., 2022; Wen et al., 2024).
 However, these successes hinge on the idealized assumption of full visibility and isolation from the background.
 This assumption rarely holds in unconstrained in-the-wild scenarios, where subjects inevitably interact with their surroundings, such as sitting on chairs, navigating behind obstacles, or being partially obscured by environmental objects.
-When standard pipelines are applied to occluded scenarios, their performance degrades catastrophically [xiang2023rendering, sun2024occfusion].
+When standard pipelines are applied to occluded scenarios, their performance degrades catastrophically (Xiang et al., 2023; Sun et al., 2024).
 As illustrated in Fig.
@@
 Consequently, enabling robust human rendering from occluded observations is imperative to bridge the gap between laboratory prototypes and real-world applications.
-To address these specific challenges, pioneering NeRF-based approaches, such as OccNeRF [xiang2023rendering] and Wild2Avatar [xiang2025rendering], explore the task of rendering humans from object-occluded monocular videos.
+To address these specific challenges, pioneering NeRF-based approaches, such as OccNeRF (Xiang et al., 2023) and Wild2Avatar (Xiang et al., 2025), explore the task of rendering humans from object-occluded monocular videos.
 By incorporating geometry priors or decoupling scene components, these methods demonstrate the feasibility of recovering occluded surfaces.
@@
 Specifically, the reliance on expensive volumetric rendering incurs prohibitive training and inference costs, effectively precluding their deployment in real-time applications.
-Recently, the advent of 3D Gaussian Splatting (3DGS) [kerbl20233d] shifted the paradigm, offering real-time rendering capabilities [fei20243d, feng2025flashgs].
+Recently, the advent of 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) shifted the paradigm, offering real-time rendering capabilities (Fei et al., 2024; Feng et al., 2025).
 To adapt 3DGS for occluded human rendering, concurrent works incorporate external priors to compensate for missing information.
-Methods relying on geometric heuristics enforce rigid rules, such as left-right symmetry [jiang2025symgaussian] or feature aggregation [ye2025occgaussian].
+Methods relying on geometric heuristics enforce rigid rules, such as left-right symmetry (Jiang et al., 2025) or feature aggregation (Ye et al., 2025).
 However, these hand-crafted spatial priors typically lack flexibility in modeling diverse human appearances.
@@
 1(f), the symmetry prior fails to recover the chest logo, erroneously copying unadorned features from the opposite side.
-Alternatively, methods leveraging generative priors utilize 2D diffusion models to hallucinate invisible regions [sun2024occfusion, lee2024guess].
+Alternatively, methods leveraging generative priors utilize 2D diffusion models to hallucinate invisible regions (Sun et al., 2024; Lee et al., 2024).
 Yet, the inherent stochasticity of frame-by-frame generation often induces severe temporal flickering.
@@
 This motivates adopting a 4D representation to consolidate fragmented observations into a unified canonical space.
-However, naive 4D Gaussian Splatting faces a supervision dilemma [wu20244d, li2024st].
+However, naive 4D Gaussian Splatting faces a supervision dilemma (Wu et al., 2024; Li et al., 2024b).
 Indiscriminately treating all pixels as valid constraints forces the deformation field to establish erroneous correspondences, compelling the human geometry to adhere to the occluder’s surface texture.
-Conversely, strictly limiting supervision to visible regions leaves the occluded parts mathematically unconstrained; without gradient feedback, the geometry in these blind spots fails to densify or maintain coherence, resulting in incomplete structures.
-Thus, the reconstruction process demands a dynamic discrimination mechanism to assess the validity of the observation, differentiating between visible human details and misleading observational noise (i.e., occluded regions).
+Conversely, a seemingly straightforward alternative is to apply hard visibility masks to exclude occluded regions from the optimization.
+Yet, strictly limiting supervision to visible regions leaves the occluded parts mathematically unconstrained; without gradient feedback, the geometry in these blind spots fails to densify or maintain coherence, resulting in incomplete structures.
+Thus, rather than relying on brittle hard masks, the reconstruction process demands a dynamic discrimination mechanism to assess the validity of the observation, differentiating between visible human details and misleading observational noise (i.e., occluded regions).
 To this end, we reformulate the task as a Maximum A Posteriori (MAP) estimation problem under heteroscedastic observation noise.
@@
 Beyond standard geometric warping, this module explicitly predicts the per-primitive aleatoric uncertainty, serving as the confidence metric for our MAP objective.
-To enable the adaptive gradient modulation, we introduce a Double Rasterization pipeline.
+To enable the adaptive gradient modulation, we introduce a Joint Rasterization pipeline.
 By rendering pixel-aligned uncertainty maps alongside color, this mechanism enforces the inverse weighting scheme, ensuring that gradients from occluded regions are effectively attenuated.
@@
 These constraints leverage the learned uncertainty to selectively propagate validity from confident regions to unreliable areas, ensuring coherent completion.
-Consequently, as illustrated in Fig.
+As illustrated in Fig.
 1(a), U-4DGS achieves high-fidelity rendering even under severe occlusion while maintaining training efficiency.
@@
 By modeling heteroscedastic observation noise, we equip the optimization with a mechanism to discriminate between valid supervision and occlusion artifacts.
-We propose U-4DGS, a framework integrating a Probabilistic Deformation Network with a Double Rasterization pipeline.
+We propose U-4DGS, a framework integrating a Probabilistic Deformation Network with a Joint Rasterization pipeline.
 This design renders pixel-aligned uncertainty maps that act as adaptive gradient modulators, effectively shielding the canonical geometry from corruption.
@@
 We demonstrate that U-4DGS achieves state-of-the-art rendering fidelity on ZJU-Mocap and OcMotion datasets, successfully recovering clean avatars from severe occlusions.
-2 Related Works
+2. Related Works
 3D Human Avatar Reconstruction.
-Traditional human reconstruction methods typically rely on dense camera arrays or depth sensors, limiting their application in in-the-wild scenarios [collet2015high, su2020robustfusion, huang2020arch, yu2021function4d].
-To address this, NeRF-based methods have enabled high-fidelity reconstruction from monocular inputs by conditioning radiance fields on SMPL priors [yu2023monohuman, te2022neural, li2024ghunerf].
+Traditional human reconstruction methods typically rely on dense camera arrays or depth sensors, limiting their application in in-the-wild scenarios (Collet et al., 2015; Su et al., 2020; Huang et al., 2020; Yu et al., 2021).
+To address this, NeRF-based methods have enabled high-fidelity reconstruction from monocular inputs by conditioning radiance fields on SMPL priors (Yu et al., 2023; Te et al., 2022; Li et al., 2024a).
 However, the reliance on volumetric ray-marching incurs prohibitive computational costs, precluding real-time applicability.
-Recently, 3D Gaussian Splatting (3DGS) [kerbl20233d] has introduced a paradigm shift with its real-time rasterization capabilities.
-State-of-the-art methods [qian20243dgs, kocabas2024hugs, hu2024gauhuman, lei2024gart, shao2024splattingavatar, zhao2025surfel] adapt this representation to dynamic humans by representing the body in a canonical space and transforming it via Linear Blend Skinning.
+Recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has introduced a paradigm shift with its real-time rasterization capabilities.
+Building upon the skeleton-conditioned canonical deformation paradigm pioneered by early implicit frameworks like Animatable NeRF (Peng et al., 2021a), SOTA methods (Qian et al., 2024; Kocabas et al., 2024; Hu et al., 2024b; Lei et al., 2024; Shao et al., 2024; Zhao et al., 2025) adapt this representation to dynamic humans by representing the body in a canonical space and transforming it via Linear Blend Skinning.
 Critically, while performant under full visibility, these standard pipelines implicitly assume valid supervision for all pixels.
@@
 Existing methodologies generally fall into three paradigms.
-Early strategies focused on scene decoupling, employing visibility priors to isolate the subject from obstacles [xiang2025rendering, xiang2023rendering]; however, these implicit frameworks often suffer from high computational overhead.
-To synthesize unobserved content, recent approaches leverage generative priors via pre-trained 2D diffusion models [sun2024occfusion, lee2024guess, fan2026inpainthuman].
+Early strategies focused on scene decoupling, employing visibility priors to isolate the subject from obstacles (Xiang et al., 2025, 2023); however, these implicit frameworks often suffer from high computational overhead.
+To synthesize unobserved content, recent approaches leverage generative priors via pre-trained 2D diffusion models (Sun et al., 2024; Lee et al., 2024; Fan et al., 2026).
 Yet, the inherent stochasticity of such processes frequently induces identity drift and temporal flickering.
-Alternatively, other 3DGS-based methods employ geometric heuristics such as symmetry constraints [ye2025occgaussian, jiang2025symgaussian], which often falter under asymmetric clothing or motions.
+Alternatively, other 3DGS-based methods employ geometric heuristics such as symmetry constraints (Ye et al., 2025; Jiang et al., 2025), which often falter under asymmetric clothing or motions.
 Distinctively, U-4DGS eschews unstable hallucinations and rigid heuristics.
@@
 Dynamic Reconstruction and Uncertainty Modeling.
-Recent strides in dynamic Gaussian Splatting have achieved high-fidelity 4D reconstruction via deformation fields [yang2024deformable, wu20244d, luiten2024dynamic, bae2024per, zhang2025mega].
+Recent strides in dynamic Gaussian Splatting have achieved high-fidelity 4D reconstruction via deformation fields (Yang et al., 2024; Wu et al., 2024; Luiten et al., 2024; Bae et al., 2024; Zhang et al., 2025).
 However, these methods predicate their success on a deterministic correspondence between canonical and observation spaces.
-While valid for clean sequences, this premise disintegrates under occlusion, where the absence of valid motion cues causes the solver to overfit to occluders [guo2025uncertainty].
-To address this, uncertainty modeling has been explored to enhance robustness against sparse views or transient noise [sunderhauf2023density, goli2024bayes, ren2024nerf].
-Notably, concurrent work like USPLAT4D [guo2025uncertainty] employs an uncertainty-aware motion graph for general dynamic scenes.
+While valid for clean sequences, this premise disintegrates under occlusion, where the absence of valid motion cues causes the solver to overfit to occluders (Guo et al., 2025).
+To address this, uncertainty modeling has been explored to enhance robustness against sparse views or transient noise (Sünderhauf et al., 2023; Goli et al., 2024; Ren et al., 2024).
+Grounded in Bayesian deep learning for aleatoric uncertainty (Kendall and Gal, 2017), prior neural rendering methods have successfully mitigated transient distractors utilizing strategies such as per-image uncertainty embeddings (Martin-Brualla et al., 2021), robust loss formulations (Sabour et al., 2023), or direct volume density parameterization (Lee et al., 2025).
+Concurrent work like USPLAT4D (Guo et al., 2025) employs an uncertainty-aware motion graph for general dynamic scenes.
 Yet, treating uncertainty solely as a smoothing weight or edge attribute proves insufficient for articulated human reconstruction, where structural integrity is paramount.
@@
 Beyond simple loss attenuation, we leverage learned confidence to actively modulate optimization dynamics: shielding canonical geometry from misleading gradients while simultaneously triggering confidence-aware regularizations to enforce physical plausibility within regions lacking reliable visual cues.
-3 Preliminaries
-3.1 3D Gaussian Splatting
-3D Gaussian Splatting (3DGS) [kerbl20233d] explicitly models the scene using a set of anisotropic 3D Gaussians.
+3. Preliminaries
+3.1. 3D Gaussian Splatting
+3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) explicitly models the scene using a set of anisotropic 3D Gaussians.
 Each primitive is parameterized by a mean , covariance , opacity , and view-dependent color (via Spherical Harmonics).
@@
 where represents the probability density of the projected 2D Gaussian at pixel , and denotes the accumulated transmittance.
-3.2 Articulated Human Gaussians
-Following recent works [hu2024gauhuman, qian20243dgs], we leverage the SMPL model [loper2015smpl] to establish a canonical coordinate system.
+3.2. Articulated Human Gaussians
+Following recent works (Hu et al., 2024b; Qian et al., 2024), we leverage the SMPL model (Loper et al., 2015) to establish a canonical coordinate system.
 SMPL provides a skinned vertex mesh parametrized by pose and shape .
@@
 This strategy ensures that coarse body movements are explicitly driven by skeletal articulation, while a learned neural field subsequently models residual non-rigid deformations (e.g., clothing dynamics).
-3.3 Problem Formulation: A Probabilistic Perspective
+3.3. Problem Formulation
 Given a monocular video sequence capturing a dynamic human, our objective is to reconstruct a high-fidelity human avatar.
@@
 4.
-4 Methodology
-4.1 Overview
+4. Methodology
+4.1. Overview
 Building upon the probabilistic formulation derived in Sec.
@@
 4.2): We introduce a network that serves a dual role: it models the temporal geometry evolution via deformation offsets and parameterizes the input-dependent aleatoric uncertainty for each Gaussian, explicitly quantifying the confidence of the temporal correspondence.
-(2) Double Rasterization (Sec.
+(2) Joint Rasterization (Sec.
 4.3): To propagate this 3D confidence to the 2D observation space, we employ a dual-branch rasterizer.
@@
 Simultaneously, to prevent geometric drift in regions lacking reliable visual cues, we enforce Confidence-Aware Regularizations that leverage the learned uncertainty to impose spatial-temporal priors.
-4.2 Probabilistic Deformation Network
+4.2. Probabilistic Deformation Network
 To instantiate the MAP framework, we require a mechanism that not only recovers time-varying geometry but also quantifies the reliability of each correspondence.
-Unlike standard deterministic fields [yang2024deformable, hu2024gauhuman] that indiscriminately fit all observations, we propose a Probabilistic Deformation Network .
+Unlike standard deterministic fields (Yang et al., 2024; Hu et al., 2024b) that indiscriminately fit all observations, we propose a Probabilistic Deformation Network .
 This module serves a dual purpose: (1) it models high-frequency non-rigid residuals beyond the SMPL prior; (2) it explicitly predicts the aleatoric uncertainty , providing the input-dependent noise scale required for our probabilistic objective.
@@
 This design allows the network to learn specific spatial-temporal failure modes, such as assigning consistently high uncertainty to body parts undergoing rapid deformation or frequent self-occlusion, regardless of their global position in the camera view.
-4.3 Double Rasterization
-To bridge the gap between the 3D aleatoric uncertainty and the 2D observation space required for our MAP objective, we propose a Double Rasterization scheme.
+[…]
+[…]
+| Standard Human Rendering | |||||||||||||
+| HumanNeRF (Weng et al., 2022) | - | 20.67 | 0.9509 | - | - | - | - | ||||||
+| GaussianAvatar (Hu et al., 2024a) | - | 18.01 | 0.9512 | 60.33 | - | - | - | ||||||
+| GauHuman (Hu et al., 2024b) | - | 21.55 | 0.9430 | 55.88 | 15.09 | 0.8525 | 107.1 | ||||||
+| Occlusion-Aware Approaches | |||||||||||||
+| OccNeRF (Xiang et al., 2023) | Scene Decoupling | 22.40 | 0.9562 | 43.01 | 15.71 | 0.8523 | 82.90 | ||||||
+| Wild2Avatar (Xiang et al., 2025) | - | - | - | 14.09 | 0.8484 | 93.31 | |||||||
+| OccGaussian (Ye et al., 2025) | Geometric Heuristics | 23.29 | 0.9482 | 41.93 | - | - | - | ||||||
+| SymGaussian (Jiang et al., 2025) | 23.22 | 0.9535 | 39.02 | - | - | - | |||||||
+| GTU (Lee et al., 2024) | Generative Priors | 22.89 | 0.9503 | 40.78 | 15.83 | 0.8437 | 83.46 | ||||||
+| OccFusion (Sun et al., 2024) | 23.96 | 0.9548 | 32.34 | 18.28 | 0.8875 | 82.42 | |||||||
+| U-4DGS (Ours) | 24.62 | 0.9606 | 31.72 | 20.11 | 0.9030 | 79.17 | |||||||
+4.3. Joint Rasterization
+To bridge the gap between the 3D aleatoric uncertainty and the 2D observation space required for our MAP objective, we propose a Joint Rasterization scheme.
 Since occlusion is an inherently view-dependent phenomenon, the reliability of a pixel is determined by the accumulated uncertainty along its corresponding optical ray.
@@
 Consequently, serves as the spatially varying noise scale, functioning as the adaptive gradient modulator in the subsequent optimization.
-4.4 Optimization via MAP Estimation
-To reconstruct high-fidelity humans from monocular videos, we jointly optimize the canonical Gaussian parameters and the deformation network weights .
+4.4. Optimization via MAP Estimation
+To reconstruct high-fidelity humans from monocular videos, we jointly optimize the Gaussian parameters and the deformation network weights .
 Following the MAP formulation introduced in Sec.
@@
 (2) Uncertainty-Weighted Temporal Inertia: To prevent jittering in the absence of visual cues, we enforce a smoothness prior on the trajectory of uncertain Gaussians.
-Considering that we train on monocular video sequences with a frame interval , we formulate this constraint as a second-order difference over the sampled timestamps:
+Considering that we train on video sequences with a frame interval , we formulate this constraint as a second-order difference over the sampled timestamps:
 where denotes the set of deformation parameters, and represents the set of valid frames with available temporal neighbors.
@@
 For high-uncertainty Gaussians, it penalizes acceleration, effectively interpolating their dynamics based on the trajectory established by confident frames, ensuring smooth motion transition across occlusion gaps.
-Note that in both and , we apply a stop-gradient operator to the uncertainty weights .
-This ensures that serves purely as an attention mechanism for the regularization strength, preventing the optimizer from trivially reducing to minimize the regularization penalty.
-Total Loss.
-The final objective function consolidates the MAP formulation, combining the uncertainty-weighted likelihood (data fidelity), the confidence-aware priors (regularization), and auxiliary perceptual constraints:
-where ensures perceptual quality using standard metrics [wang2004image, zhang2018unreasonable].
+In both and , we apply a stop-gradient operator to the uncertainty weights to prevent degenerate optimization.
+Because scales the regularization penalties, unconstrained backpropagation would encourage the network to trivially drive simply to minimize these losses, invalidating its role in modeling heteroscedastic observation noise.
+By decoupling the gradient flow, ensures that the aleatoric uncertainty is learned exclusively from the photometric likelihood, functioning purely as a non-differentiable attention mask for the spatial-temporal constraints.
+Total Loss. The final objective function combines the uncertainty-weighted likelihood (data fidelity), the confidence-aware priors (regularization), and auxiliary perceptual constraints:
+where ensures perceptual quality using standard metrics (Wang et al., 2004; Zhang et al., 2018).
 Additionally, imposes a global silhouette constraint, aligning the accumulated opacity with the projected SMPL mask to guide coarse geometry.
-The hyperparameters balance the contribution of each term.
+balances the contribution of each term.
 […]
-[…]
-| Standard Human Rendering | |||||||||||||
-| HumanNeRF [weng2022humannerf] | - | 20.67 | 0.9509 | - | - | - | - | ||||||
-| GaussianAvatar [hu2024gaussianavatar] | - | 18.01 | 0.9512 | 60.33 | - | - | - | ||||||
-| GauHuman [hu2024gauhuman] | - | 21.55 | 0.9430 | 55.88 | 15.09 | 0.8525 | 107.1 | ||||||
-| Occlusion-Aware Approaches | |||||||||||||
-| OccNeRF [xiang2023rendering] | Scene Decoupling | 22.40 | 0.9562 | 43.01 | 15.71 | 0.8523 | 82.90 | ||||||
-| Wild2Avatar [xiang2025rendering] | - | - | - | 14.09 | 0.8484 | 93.31 | |||||||
-| OccGaussian [ye2025occgaussian] | Geometric Heuristics | 23.29 | 0.9482 | 41.93 | - | - | - | ||||||
-| SymGaussian [jiang2025symgaussian] | 23.22 | 0.9535 | 39.02 | - | - | - | |||||||
-| GTU [lee2024guess] | Generative Priors | 22.89 | 0.9503 | 40.78 | 15.83 | 0.8437 | 83.46 | ||||||
-| OccFusion [sun2024occfusion] | 23.96 | 0.9548 | 32.34 | 18.28 | 0.8875 | 82.42 | |||||||
-| U-4DGS (Ours) | 24.62 | 0.9606 | 31.72 | 20.11 | 0.9030 | 79.17 | |||||||
-5 Experiments
-5.1 Experimental Setup
+| A | Baseline | 22.85 | 0.9490 | 45.12 |
+| B | + Uncert. Modeling | 24.15 | 0.9575 | 35.20 |
+| C | + Spa. Con. () | 24.48 | 0.9592 | 32.50 |
+| D | + Temp. Con. () | 24.62 | 0.9606 | 31.72 |
+5. Experiments
+5.1. Experimental Setup
 Datasets.
 We evaluate U-4DGS on two standard benchmarks representing both synthetic and real-world occlusion scenarios.
-First, we utilize the ZJU-MoCap dataset [peng2021neural], which contains multi-view sequences of 6 subjects.
-Following the protocol of OccNeRF [xiang2023rendering], we simulate occlusions by masking the central 50% of the human region in the first 80% of the frames.
+First, we utilize the ZJU-MoCap dataset (Peng et al., 2021b), which contains multi-view sequences of 6 subjects.
+Following the protocol of OccNeRF (Xiang et al., 2023), we simulate occlusions by masking the central 50% of the human region in the first 80% of the frames.
 We adopt a monocular setting for training, using only Camera 1 with 100 frames subsampled at an interval of 5, while the remaining 22 cameras are reserved for novel view evaluation.
-Second, to assess robustness in natural environments, we employ the OcMotion dataset [huang2022occluded], which features varying degrees of real-world human-object interactions.
-Consistent with prior works [xiang2025rendering, sun2024occfusion], we select 6 representative sequences and train on sparse sub-sequences of 50 frames to challenge the model’s capability under limited observations.
+Second, to assess robustness in natural environments, we employ the OcMotion dataset (Huang et al., 2022), which features varying degrees of real-world human-object interactions.
+Consistent with prior works (Xiang et al., 2025; Sun et al., 2024), we select 6 representative sequences and train on sparse sub-sequences of 50 frames to challenge the model’s capability under limited observations.
 Baselines.
 We benchmark our method against comprehensive state-of-the-art approaches, categorized into two distinct groups: (1) Standard Human Rendering.
-To demonstrate the impact of occlusion on conventional pipelines, we evaluate HumanNeRF [weng2022humannerf], GaussianAvatar [hu2024gaussianavatar], and GauHuman [hu2024gauhuman].
+To demonstrate the impact of occlusion on conventional pipelines, we evaluate HumanNeRF (Weng et al., 2022), GaussianAvatar (Hu et al., 2024a), and GauHuman (Hu et al., 2024b).
 These methods are trained directly on the occluded sequences without specific handling mechanisms.
 (2) Occlusion-Aware Approaches.
-We compare against representative methods covering three mainstream technical paradigms: a) Scene Decoupling: OccNeRF [xiang2023rendering] and Wild2Avatar [xiang2025rendering], which utilize implicit fields or layer-wise parameterization to separate the human subject from environmental obstacles.
-b) Geometric Heuristics: OccGaussian [ye2025occgaussian] and SymGaussian [jiang2025symgaussian], which rely on rigid priors such as feature aggregation or left-right symmetry to infer missing geometry.
-c) Generative Priors: OccFusion [sun2024occfusion] and GTU [lee2024guess], which leverage 2D diffusion models to hallucinate the unobserved regions.
+We compare against representative methods covering three mainstream technical paradigms: a) Scene Decoupling: OccNeRF (Xiang et al., 2023) and Wild2Avatar (Xiang et al., 2025), which utilize implicit fields or layer-wise parameterization to separate the human subject from environmental obstacles.
+b) Geometric Heuristics: OccGaussian (Ye et al., 2025) and SymGaussian (Jiang et al., 2025), which rely on rigid priors such as feature aggregation or left-right symmetry to infer missing geometry.
+c) Generative Priors: OccFusion (Sun et al., 2024) and GTU (Lee et al., 2024), which leverage 2D diffusion models to hallucinate the unobserved regions.
 For a fair comparison, all methods are evaluated using identical segmentation masks and SMPL pose priors.
-Implementation Details: Our framework is implemented in PyTorch [paszke2019pytorch] and trained on a single NVIDIA A100 PCIe GPU.
+Implementation Details: Our framework is implemented in PyTorch (Paszke et al., 2019) and trained on a single NVIDIA A100 PCIe GPU.
 The Probabilistic Deformation Network is instantiated as an 8-layer MLP (hidden dimension 256) with a skip connection at the 4-th layer.
 To capture high-frequency spatial-temporal dynamics, we apply sinusoidal positional encoding to input coordinates and time with frequency bands and , respectively.
-We optimize the model for 10k iterations using the Adam optimizer [kingma2014adam].
-Learning rates for canonical Gaussian attributes follow the schedule of GauHuman [hu2024gauhuman], while the deformation network utilizes a rate of .
+We optimize the model for 10k iterations using the Adam optimizer (Kingma and Ba, 2015).
+Learning rates for canonical Gaussian attributes follow the schedule of GauHuman (Hu et al., 2024b), while the deformation network utilizes a rate of .
 Crucially, to preclude trivial solutions (where the model simply predicts infinite uncertainty to minimize the NLL), we adopt a progressive training strategy.
@@
 The regularization weights are set as: , , and .
-5.2 Comparisons with State-of-the-Arts
+5.2. Comparisons with State-of-the-Arts
 Quantitative Results.
@@
 Notably, it surpasses the leading generative baseline, OccFusion (23.96 dB), validating that our uncertainty-guided aggregation yields superior fidelity without relying on potentially unstable diffusion priors.
-Crucially, on the real-world OcMotion dataset, our framework demonstrates exceptional robustness, outperforming the runner-up by a substantial margin of 1.83 dB.
+On the real-world OcMotion dataset, U-4DGS demonstrates exceptional robustness.
+Since the ground-truth of human appearance for occluded regions in real-world scenarios is unavailable, we calculate the metrics on visible pixels only.
+U-4DGS outperforms the runner-up by a substantial margin of 1.83 dB.
 Unlike standard rendering or scene decoupling methods that degrade under occlusions, U-4DGS effectively leverages Confidence-Aware Regularizations to shield the canonical geometry, proving that modeling heteroscedastic noise is decisive for handling unconstrained interactions.
@@
 3 visualizes the rendering quality.
-Standard methods like GauHuman (GH) [hu2024gauhuman] lack occlusion awareness, erroneously fusing environmental obstacles (e.g., bars or boxes) directly onto the human body.
-While generative approaches like OccFusion (OF) [sun2024occfusion] can remove occluders, they suffer from inherent stochasticity, often yielding over-smoothed textures or hallucinations inconsistent with the subject’s identity.
+Standard methods like GauHuman (GH) (Hu et al., 2024b) lack occlusion awareness, erroneously fusing environmental obstacles (e.g., bars or boxes) directly onto the human body.
+While generative approaches like OccFusion (OF) (Sun et al., 2024) can remove occluders, they suffer from inherent stochasticity, often yielding over-smoothed textures or hallucinations inconsistent with the subject’s identity.
 In contrast, U-4DGS achieves superior fidelity.
@@
 By strictly aggregating valid temporal information, our method faithfully restores clean, sharp, and physically consistent avatars, even under severe occlusions.
-5.3 Ablation and Analysis
+5.3. Ablation and Analysis
 Ablation Studies.
@@
 (1) Exp.
-A (Baseline): We establish a deterministic baseline OccGauHuman, structurally inheriting the pipeline from GauHuman [hu2024gauhuman].
+A (Baseline): We establish a deterministic baseline OccGauHuman, structurally inheriting the pipeline from GauHuman (Hu et al., 2024b).
 It employs a standard deformation field optimized via loss, effectively assuming homoscedastic observation noise.
@@
 (2) Exp.
-B (+ Uncertainty Modeling): Replacing the deterministic field with our Probabilistic Deformation Network and enabling Double Rasterization yields a substantial quality leap.
+B (+ Uncertainty Modeling): Replacing the deterministic field with our Probabilistic Deformation Network and enabling Joint Rasterization yields a substantial quality leap.
 This step introduces the Likelihood term of our MAP objective.
@@
 This temporal prior ensures trajectory smoothness and further refines surface details, achieving the most faithful reconstruction with high physical plausibility.
-[…]
-| A | Baseline | 22.85 | 0.9490 | 45.12 |
-| B | + Uncertainty Modeling | 24.15 | 0.9575 | 35.20 |
-| C | + Spatial Constraint () | 24.48 | 0.9592 | 32.50 |
-| D | + Temporal Constraint () | 24.62 | 0.9606 | 31.72 |
 Visualization of Uncertainty Maps.
@@
 5.
-In our MAP framework, represents the estimated scale of observation noise, tasked with autonomously distinguishing between valid signals and occlusion artifacts.
+In our MAP framework, denotes the estimated scale of observation noise and is tasked with autonomously distinguishing valid signals from occlusion artifacts.
 As evident in Fig.
@@
 7), it effectively attenuates gradients in these regions lacking reliable visual cues, preventing the canonical geometry from adhering to environmental occlusions.
-Temporal Stability Analysis.
-To validate the temporal coherence of our framework, we conduct a longitudinal comparison against the generative baseline GTU [lee2024guess] across a long sequence (Fig.
-6).
-As observed in the top row, GTU suffers from severe identity drift due to the inherent stochasticity of its diffusion priors.
-Since the method hallucinates missing regions frame-by-frame without strict temporal coupling, the synthesized textures fluctuate drastically.
-This instability becomes particularly pronounced in later frames (e.g., to ), where the clothing consistency collapses, degenerating into severe dark artifacts and incoherent noise on the subject’s back.
-In contrast, U-4DGS achieves rigorous stability.
-By anchoring appearance in a unified canonical space and enforcing Uncertainty-Weighted Temporal Inertia, our method effectively constrains the optimization trajectory.
-This ensures that the reconstructed human maintains physical consistency and sharp details throughout the sequence, eliminating the texture flickering observed in generative approaches.
-6 Conclusion
+[…]
+| OccNeRF | 20 h | 0.20 |
+| OccGaussian | 6 min | 169 |
+| OccFusion | 10 min | 151 |
+| U-4DGS (Ours) | 7.5 min | 155 |
+Efficiency Analysis.
+We present a quantitative comparison of training time and rendering speed (FPS) in Table 3.
+Implicit methods such as OccNeRF are severely bottlenecked by volumetric ray-marching, resulting in prohibitive training times and non-real-time inference.
+Generative prior-based approaches like OccFusion achieve real-time rendering but require relatively longer training (10 minutes) due to the integration of diffusion models.
+In contrast, U-4DGS preserves the efficiency inherent to 3D Gaussian Splatting, achieving an interactive rendering speed of 155 FPS.
+Furthermore, the introduction of the Probabilistic Deformation Network and the joint rasterization pipeline introduces minimal computational overhead compared to heuristic-based methods (e.g., OccGaussian), requiring only 7.5 minutes for training.
+6. Conclusion
 In this paper, we present U-4DGS, a robust framework that reformulates monocular human rendering as an MAP estimation problem to handle severe environmental occlusion.
 Departing from paradigms that rely on stochastic hallucinations or rigid priors, we explicitly model the heteroscedastic nature of observation noise.
-By integrating a Probabilistic Deformation Network with a Double Rasterization pipeline, our method establishes an adaptive gradient modulator, enabling the optimization to autonomously distinguish between valid geometric cues and occlusion artifacts.
+By integrating a Probabilistic Deformation Network with a Joint Rasterization pipeline, our method establishes an adaptive gradient modulator, enabling the optimization to autonomously distinguish between valid geometric cues and occlusion artifacts.
 Furthermore, we leverage the learned uncertainty to drive Confidence-Aware Regularizations, preventing geometric drift and ensuring structural integrity in regions lacking reliable visual cues.